# Self-Supervised Monocular 4D Scene Reconstruction for Egocentric Videos

### v3

@@ Abstract
[…]
@@ 1 Introduction
Figure 2: The overview of EgoMono4D and our self-supervised training framework. The model first simultaneously predicts camera intrinsic, video depth, and confidence maps (for camera pose estimation). Camera poses are then calculated by aligning unprojected pointclouds from different frames with confidence maps. The final dense pointclouds sequence reconstruction is assembled using all the predicted variables. We train our model purely on unlabeled egocentric video datasets, with both self-supervised photometric loss for depth alignment and regularization loss for training stablization.
@@ 5.1 Datasets
Figure 4: The visualization of the dense pointcloud sequence reconstruction by EgoMono4D demonstrates its ability to effectively recover both the overall scene structure and dynamic motion elements to a significant extent. For additional visualizations, please refer to Appendix G. We also provide a qualitative comparison and visualization with baseline methods in Appendix H. Video and interactable visualizations can be found in project website.
Figure 5: Inference speed comparison of different methods.
@@ 5.3 Baseline
We compare our model to previous methods that (1) provide dense 4D reconstruction, (2) have nearly linear time complexity with respect to the number of frames, enabling large-scale reconstruction, and (3) demonstrate zero-shot generalization. DS+UniDepth (DROID-SLAM [77] + UniDepth [56]) combines depth estimation with learning-based RGBD visual odometry. MFVR+UniDepth (MapFreeVR [1] + UniDepth [56]) first estimates image depth, then computes camera poses with correspondence estimation (optical flow [92] in our setting) and depth-alignment. Additionally, we integrate HOI masks from EgoHOS[100] to filter out the HOI region for alignment. This baseline could be viewed as a modularized and no-training version of EgoMono4D. DUSt3R [86] supports end-to-end reconstruction. We use the ”swin” mode from the original implementation to achieve $\mathcal{O}(T)$ inference. MonSt3R [99] is a finetuned version of DUSt3R on synthetic dynamic datasets, aimed at improving reconstruction performance in dynamic areas. Lastly, Align3R [45] further merge depth prior during finetuning to enhance the geometry estimation ability.
@@ 6 Result
Here we present the evaluation and ablation results of EgoMono4D. Detailed definitions of all metrics can be found in Appendix C.
@@ 6.1 Dense Pointclouds Sequence Reconstruction
Results  The quantitative evaluation results are presented in Table 1, with visualizations shown in Figure 4. For additional visualizations, refer to Appendix G. We also provide a qualitative comparison and visualization with baseline methods in Appendix H. Video and interactable visualization can be found in project website .
EgomMono4D demonstrates superior performance across all evaluated methods. Notably, on the challenging POV-Surgery dataset [85], which contains complex surgical scenes with unrealistic textures and intricate actions, our model outperforms others by 10–20% in terms of F-score. DS+UniDepth also shows commendable performance across several metrics. However, qualitative visualizations (Appendix H) reveal that it struggles to align the static portions of egocentric scenes.
The modular version of our model, MFVR+UniDepth, performs significantly worse across most metrics. This can be attributed to inconsistencies between the different modules. In contrast, our end-to-end self-supervised training (described in Section 4.3) ensures that the modules align through back-propagation of the 4D supervision loss, leading to improved 4D reconstruction accuracy.
Due to the limited availability of labeled egocentric data for training, DUSt3R, MonSt3R and Align3R exhibit suboptimal performance on egocentric videos. Furthermore, fine-tuning MonSt3R on small-scale dynamic datasets actually results in performance degradation compared to DUSt3R, primarily due to domain gap and overfitting. By introducinging depth prior, Align3R alleviates this issue, but still fails to get precise monocular 4D estimation. These results highlight the challenges that supervised learning methods face in label-scarce scenarios. In contrast, our self-supervised approach performs effectively on unlabeled egocentric videos. It is important to note that this does not imply that supervised methods are inherently inferior to self-supervised ones. In domains with abundant labeled data, supervised methods may offer advantages, as demonstrated in depth estimation tasks [61].
@@ 6.2 Long-term 3D Scene Flow Recovery
6.3 Inference Speed
We evaluate the inference speed of all models (for 40 frames video). All measurements are conducted on a single NVIDIA GeForce 3090 GPU with a batch size of 1. Results are shown in Figure 5. Our model achieves the fastest speed, except for its modularized version (MFVR+UniDepth).
@@ 6.4 Ablation Study
We also conduct an ablation on the hyperparameters for model inference, including the window size $N_{w}$ and window overlap size $N_{o}$, which are detailed in Appendix E. For the window size $N_{w}$, maintaining consistency between training and inference is critical, likely because the video adaptor is specifically trained to fuse 4 frames. Regarding the overlap size $N_{o}$, the model performs comparably for $N_{o}=1,2,3$. Therefore, we select $N_{o}=1$ to maximize inference speed.
@@ 7 Conclusion
We present EgoMono4D, an self-supervised model for 4D reconstruction of egocentric videos, trained solely on large-scale unlabeled data. By aligning video depth with 4D constraints, it achieves promising zero-shot results in dense, generalizable scene reconstruction. Additional visualization, details, discussion, limitations and future direction could be found in appendix.
@@
  • E. Ablation of Inference Strategy
  • F. Depth and Camera Results
  • G. More Reconstruction Results Visualization
  • H. Qualitative Comparison with Baseline
  • I. Limitations and Future Directions
For video and interactable visualization, please refer to the project website.
@@ Appendix D Long-term 3D Scene Flow Recovery Results
Appendix E Ablation on Inference Strategy
@@
Appendix F Depth and Camera Results
@@
Appendix G More Reconstruction Results Visualization
@@
Appendix H Qualitative Comparison with Baseline
@@
Appendix I Limitations and Future Directions

### v4

@@ Abstract
footnotetext: † Work done during the internship at Shanghai Qi Zhi Institute. ‡ The corresponding author.
@@ 1 Introduction
Figure 2: The overview of EgoMono4D and our self-supervised training framework. The model first simultaneously predicts camera intrinsic, video depth, and confidence maps (for camera pose estimation). Camera poses are then calculated by aligning unprojected pointclouds from different frames with confidence maps. The final dense pointclouds sequence reconstruction is assembled using all the predicted variables. We train our model purely on unlabeled egocentric video datasets, with both self-supervised photometric loss for depth alignment and regularization loss for training stabilization.
@@ 5.1 Datasets
Figure 4: The visualization of the dense pointcloud sequence reconstruction by EgoMono4D demonstrates its ability to effectively recover both the overall scene structure and dynamic motion elements to a significant extent. For additional visualizations, please refer to Appendix H. We also provide a qualitative comparison and visualization with baseline methods in Appendix I. Video visualizations and interactive demonstration could be found in https://egomono4d.github.io/.
@@ 5.3 Baseline
We compare our model to previous methods that (1) provide dense 4D reconstruction, (2) have nearly linear time complexity with respect to the number of frames, enabling large-scale reconstruction, and (3) demonstrate zero-shot generalization. Modularized Version (MapFreeVR [1] + UniDepth [56]) first estimates image depth, then computes camera poses with correspondence estimation (optical flow [92] in our setting) and depth-alignment. Additionally, we integrate HOI masks from EgoHOS[100] to filter out the HOI region for alignment. This baseline could be viewed as a modularized and no-training version of EgoMono4D. DS+UniDepth (DROID-SLAM [77] + UniDepth [56]) combines depth estimation with learning-based RGBD visual odometry. DUSt3R [86] supports end-to-end reconstruction. We use the "swin" mode from the original implementation to achieve $\mathcal{O}(T)$ inference. MonSt3R [99] is a finetuned version of DUSt3R on synthetic dynamic datasets, aimed at improving reconstruction performance in dynamic areas. Align3R [45] further merges depth prior during finetuning to enhance the geometry estimation ability. Lastly, CUT3R [84] incorporates memory buffers for temporal information merging to improve reconstruction consistency.
@@ 6 Result
Here we present the evaluation and ablation results of EgoMono4D. Detailed definitions of all metrics can be found in Appendix C. We include the speed measurement and comparison of all models in Appendix E. We also include the evaluation results of depth and camera poses estimation in Appendix G.
@@ 6.1 Dense Pointclouds Sequence Reconstruction
Results  The quantitative evaluation results are presented in Table 1, with visualizations shown in Figure 4. For additional visualizations, refer to Appendix H. We also provide a qualitative comparison and visualization with baseline methods in Appendix I. Video and interactive demonstration visualizations can be found in project website.
EgoMono4D demonstrates superior performance across all evaluated methods. Notably, on the challenging POV-Surgery dataset [85], which contains complex surgical scenes with unrealistic textures and intricate actions, our model outperforms others by 10–20% in terms of F-score. DS+UniDepth also shows commendable performance across several metrics. However, qualitative visualizations (Appendix I) reveal that it struggles to align the static portions of egocentric scenes.
The modularized version of our model performs significantly worse across most metrics. This can be attributed to inconsistencies between the different modules. In contrast, our end-to-end self-supervised training (described in Section 4.3) ensures that the modules align through back-propagation of the 4D supervision loss, leading to improved 4D reconstruction accuracy.
Due to the limited availability of labeled egocentric data for training, DUSt3R, MonSt3R, Align3R and CUT3R exhibit suboptimal performance on egocentric videos. Furthermore, fine-tuning MonSt3R on small-scale dynamic datasets actually results in performance degradation compared to DUSt3R, primarily due to domain gap and overfitting. By introducing depth prior, Align3R alleviates this issue, but still fails to get precise monocular 4D estimation. These results highlight the challenges that supervised learning methods face in label-scarce scenarios. In contrast, our self-supervised approach performs effectively on unlabeled egocentric videos. It is important to note that this does not imply that supervised methods are inherently inferior to self-supervised ones. In domains with abundant labeled data, supervised methods may offer advantages, as demonstrated in depth estimation tasks [61].
@@ 6.3 Ablation Study
We also conduct an ablation on the hyperparameters for model inference, including the window size $N_{w}$ and window overlap size $N_{o}$, which are detailed in Appendix F. For the window size $N_{w}$, maintaining consistency between training and inference is critical, likely because the video adaptor is specifically trained to fuse 4 frames. Regarding the overlap size $N_{o}$, the model performs comparably for $N_{o}=1,2,3$. Therefore, we select $N_{o}=1$ to maximize inference speed.
@@ 7 Conclusion
We present EgoMono4D, an self-supervised model for 4D reconstruction of egocentric videos, trained solely on large-scale unlabeled data. By aligning video depth with 4D constraints, it achieves promising zero-shot results in dense, generalizable scene reconstruction. Additional discussion of limitations and future directions of EgoMono4D could be found in Appendix J.
8 Acknowledgment
This work is supported by the National Key R&D Program of China (2022ZD0161700), National Natural Science Foundation of China (62176135), Shanghai Qi Zhi Institute Innovation Program SQZ202306 and the Tsinghua University Dushi Program, the grant of National Natural Science Foundation of China (NSFC) 12201341.
@@ 8 Acknowledgment
  • E. Results of Model Inference Speed.
  • F. Ablation of Inference Strategy
  • G. Depth and Camera Results
  • H. More Reconstruction Results Visualization
  • I. Qualitative Comparison with Baseline
  • J. Limitations and Future Directions
For video visualization and online interactive demonstration, please refer to https://egomono4d.github.io/.
@@ Appendix D Long-term 3D Scene Flow Recovery Results
Appendix E Inference Speed Results
Figure 10: Inference speed comparison of different methods.
We evaluate the inference speed of different models (for 40 frames video). All measurements are conducted on a single NVIDIA GeForce 3090 GPU with a batch size of 1. Results are shown in Figure 10. Our model achieves 0.218 secs/frame speed, which is only slower than its modularized version [1] (0.215 secs/frame) and CUT3R [84] (0.115 sec/frame). However, the speed difference is very small, while our model achieves better reconstruction quality with large advantage.
Appendix F Ablation on Inference Strategy
@@
Appendix G Depth and Camera Results
@@
We believe this is due to two main reasons: (1) Our method is primarily optimized for 4D pointclouds reconstruction, rather than single geometry variable estimation. While it may not achieve the highest accuracy for individual variables, it ensures greater consistency and alignment among them. (2) Compared with supervised methods, self-supervised approach does not show a clear advantage when focusing solely on single-variable estimation, as indicated by previous work [72]. This is because there are more datasets available for single-variable estimation. We believe that using more advanced pretrained models and training on 0larger datasets may address this in the future [72, 73].
Appendix H More Reconstruction Results Visualization
@@
Appendix I Qualitative Comparison with Baseline
@@
Appendix J Limitations and Future Directions